OpenStudy (anonymous):

find the value of x in the proportion, plz show your work so i can understand! x/ x-12 = 9/5

OpenStudy (anonymous):

multiply both sides of the equation ( to get rid of nasty fraction): you multiply by 5(x - 12): 5x = 9(x-12) 5x = 9x - 108 subtract 5x fro both sides: 0 = 4x - 108 add 108 to both sides 108 = 4x x = 108/4 = 27